# Parametricity via Cohesion

### v1

@@ Abstract
Parametricity is a key metatheoretic property of type systems, which implies strong uniformity & modularity properties of the structure of types within systems possessing it. In recent years, various systems of dependent type theory have emerged with the aim of expressing such parametric reasoning in their internal logic, toward the end of solving various problems arising from the complexity of higher-dimensional coherence conditions in type theory. This paper presents a first step toward the unification, simplification, and extension of these various methods for internalizing parametricity. Specifically, I argue that there is an essentially modal aspect of parametricity, which is intimately connected with the category-theoretic concept of cohesion. On this basis, I describe a general categorical semantics for modal parametricity, develop a corresponding framework of axioms (with computational interpretations) in dependent type theory that can be used to internally represent and reason about such parametricity, and show this in practice by implementing these axioms in Agda and using them to verify parametricity theorems therein. I then demonstrate the utility of these axioms in managing the complexity of higher-dimensional coherence by deriving induction principles for higher inductive types, and in closing, I sketch the outlines of a more general synthetic theory of parametricity, with applications in domains ranging from homotopy type theory to the analysis of program modules.
@@
1 The past, present & future of parametricity in type theory
@@
Reynolds [Rey83] began his seminal introduction of the concept of parametricity with the declaration that “type structure is a syntactic discipline for enforcing levels of abstraction.” In the ensuing decades, this idea of Reynolds’ has been overwhelmingly vindicated by the success of type systems in achieving modularity and abstraction in domains ranging from programming to interactive theorem proving. Yet the fate of Reynolds’ particular strategy for formalizing this idea is rather more ambiguous.
@@
Reynolds’ original analysis of parametricity targeted the polymorphic $\lambda$-calculus (aka System F). Intuitively, polymorphic programs in System F cannot inspect the types over which they are defined and so must behave essentially the same for all types at which they are instantiated. To make this intuitive idea precise, Reynolds posed an ingenious solution in terms of logical relations, whereby every System F type is equipped with a suitable binary relation on its inhabitants, such that all terms constructible in System F must preserve the relations defined on their component types. On this basis, Reynolds was able to establish many significant properties of the abstraction afforded by System F’s type structure, e.g: all closed terms of type $\forall X.X\to X$ are extensionally equivalent to the identity function.
Reynolds’ results have subsequently been extended to many type systems. However, as the complexity and expressivity of these systems increases, so too does the difficulty of defining parametricity relations for their types, and proving that the inhabitants of these types respect their corresponding relations. Moreover, the abstract, mathematical patterns underlying the definitions of such relations have not always been clear. The problem essentially is that the usual theory of parametricity relations for type systems is “analytic,” i.e. defined in terms of the particular constructs afforded by those systems, when what would be more desirable is a “synthetic” theory that boils the requirements for parametricity down to a few axioms.
@@
The need for such an axiomatic specification of parametricity is made all the more apparent by recent developments in Homotopy Type Theory (HoTT) and related fields, wherein dependent type theories capable of proving parametricity theorems internally for their own type structures have been developed and applied by various authors to solve some major open problems in these fields (c.f. [CH20], [KS24]). Each such type theory has posed its own solution to the problem of internalizing parametricity, and although some commonalities exist between these systems, there is as yet no one framework that subsumes them all. In particular, the majority of these type theories have targeted only specific semantic models (usually some appropriately-chosen presheaf categories) rather than an axiomatically-defined class of models, and in this sense the analysis of parametricity offered by these type theories remains analytic, rather than synthetic. This in particular limits any insight into whether and how these approaches to internal parametricity may be related to one another, generalized, or simplified.
@@
As a first step toward the unification, simplification, and generalization of these approaches to internal parametricity, I propose an analysis of parametricity in terms of the category-theoretic concept of cohesion. Cohesion here refers to a particular relation between categories whereby one is equipped with an adjoint string of modalities that together make its objects behave like abstract spaces, whose points are bound together by some sort of cohesion that serves to constrain the maps that exist between these spaces. Such a notion of cohesion is reminiscent of the informal idea behind Reynolds’ formulation of parametricity outlined above, particularly if one interprets the cohesion that binds types together as the structure of relations that inhere between them. Moreover, by inspection of the categorical models of extant type theories for internal parametricity, along with classical models of parametricity, one finds that many (if not quite all) of these exhibit cohesion, in this sense. The main contribution of this paper is thus twofold: 1) to show that this basic setup of cohesion is essentially all that is needed to recover classical parametricity results internally in dependent type theory, and 2) to show how this axiomatic framework can be applied in solving problems arising from the complexity of coherence conditions in HoTT – specifically the problem of deriving induction principles for higher inductive types from their recursors. As a further illustration and verification of this idea, this paper is also a literate Agda document wherein the axioms for and theorems resulting from such parametricity via cohesion have been fully formalized and checked for validity. [^1] The full source of this file can be found at:
https://github.com/cbaberle/Parametricity-via-Cohesion/blob/main/parametricity-via-cohesion.lagda.md
{-# OPTIONS --rewriting --cohesion --flat-split --without-K #-}module parametricity-via-cohesion where
@@ 2 Cohesion & Parametricity
where $\mathcal{E,S}$ are both topoi, $\Delta,\nabla$ are both fully faithful, and $\Pi$ preserves finite products. Given such an arrangement, we think of the objects of $\mathcal{E}$ as spaces and those of $\mathcal{S}$ as sets (even if $\mathcal{S}$ is not the category of sets), where $\Gamma$ is the functor that sends a space to its set of points, $\Delta$ sends a set to the corresponding discrete space, $\nabla$ sends a set to the corresponding codiscrete space, and $\Pi$ sends a space to its set of connected components. These in turn induce a string of adjoint modalities on $\mathcal{E}$:
@@
A concrete example of cohesion comes from the category of reflexive graphs $\mathbf{RGph}$, which is cohesive over the category of sets $\mathbf{Set}$ [Law05]. Here, $\Gamma$ is the functor that sends a reflexive graph to its set of vertices, $\Delta$ sends a set $V$ to the “discrete” reflexive graph on $V$ whose only edges are self-loops, $\nabla$ sends $V$ to the “codiscrete” graph where there is a unique edge between any pair of vertices, and $\Pi$ sends a reflexive graph to its set of (weakly) connected components. It is worth noting, at this point, that many classical models of parametricity (e.g. [Atk12]) are based upon semantic interpretations of type structure in terms of reflexive graphs. This, I wish to argue, is no accident, and the key property of reflexive graphs underlying such interpretations is precisely their cohesive structure. More generally, for any base topos $\mathcal{S}$, we may construct its corresponding topos $\mathbf{RGph}(\mathcal{S})$ of internal reflexive graphs, which will similarly be cohesive over $\mathcal{S}$, so we can in fact use the language of such internal reflexive graphs to derive parametricity results for any topos (or indeed, any $\infty$-topos, as described below).
In fact, this same setup of cohesion is interpretable, mutatis mutandis, in the case where $\mathcal{E,S}$ are not topoi, but rather $\infty$-topoi, i.e. models of homotopy type theory [Shu18]. This allows us to use the language of homotopy type theory – suitably extended with constructs for the above-described modalities (the $\flat$ modality in particular, which, for technical reasons, cannot be axiomatized directly in ordinary HoTT) – to work synthetically with the structure of such a cohesive $\infty$-topos. For present purposes, we accomplish this by working in Agda with the --cohesion and --flat-split flags enabled, along with --without-K, which ensures compatibility with the treatment of propositional equality in HoTT.
@@
    J⁻¹ :  {ℓ κ} {: Set ℓ} {: A}           (: (: A)  a ≡ b  Set κ)           {: A}  (: a ≡ b)  B b p  B a refl    J⁻¹ B refl b = b    transp⁻¹ :  {ℓ κ} {: Set ℓ} {a b : A}                (: A  Set κ)  (a ≡ b)  B b  B a    transp⁻¹ B p b = J⁻¹  a _  B a) p b
@@
The notion of contractibility then expresses the idea that a type is essentially uniquely inhabited.
@@
The reader familiar with HoTT may note that, so far, we have not included anything relating to the univalence axiom – arguably the characteristic axiom of HoTT. In fact, this is by design, as a goal of the current formalization is to assume only axioms that can be given straightforward computational interpretations that preserve the property that every closed term of the ambient type theory evaluates to a canonical normal form (canonicity), so that these axioms give a constructive and computationally sound interpretation of parametricity. While the univalence axiom can be given a computational interpretation compatible with canonicity, as in Cubical Type Theory [CHS22], doing so is decidedly not a straightforward matter. Moreover, it turns out that the univalence axiom is largely unneeded in what follows, save for demonstrating admissibility of some additional axioms which permit more straightforward computational interpretations that are (conjecturally) compatible with canonicity. I thus shall have need for univalence only as a metatheoretic assumption. In this setting, univalence allows us to convert equivalences between types into identifications of those types, which may then be transported over accordingly.
@@
Of principal importance here is the $\flat$ modality, which (intuitively) takes a type $A$ to its corresponding discretization $\flat A$, wherein all cohesion between points has been eliminated. However, in order for this operation to be well-behaved, $A$ must not depend upon any variables whose types are not themselves discrete, in this sense. To solve this problem, the --cohesion flag introduces a new form of variable binding @♭ x : X, which metatheoretically asserts that $x$ is an element of the discretization of $X$, such that $X$ may only depend upon variables bound with @♭. In this case, we say that $x$ is a crisp element of $X$.
With this notion in hand, we can define the $\flat$ modality as an operation on crisp types:
@@
As expected, the $\flat$ modality is a comonad with the following counit operation $\epsilon$:
    ε : {@♭ l : Level} {@♭ A : Set l}  ♭ A  A    ε (con x) = x
@@
Beyond such notions of discreteness, etc., what more is required for the sake of parametricity is some way of detecting when the elements of a given type are related, or somehow bound together, by the cohesive structure of that type.
@@
  • Strictly bipointed means that $I$ is equipped with a choice of two elements $i_{0},i_{1}:I$, such that the proposition $(i_{0}=i_{1})\to\bot$ (i.e. $i_{0}\neq i_{1}$) holds in the internal language of $\mathcal{E}$.
@@
We could also, in principle, directly postulate the strict bipointedness of $I$, as an axiom having the form i0 ≡ i1 → ⊥. However, this is in fact unnecessary, as this axiom will instead follow from an equivalent formulation introduced in a following subsection.
@@
A type $A$ is connected if and only if, for every discrete type $B$, any function $A\to B$ is essentially constant, in the sense of factoring through a contractible type.
@@
[…]
@@
In order to conveniently express this property in Agda, it shall therefore be prudent first to introduce some additional constructs for ergonomically handling paths, analogous to the definition of path types in Cubical Type Theory (as in e.g. [CCHM15], [${}^{+}$21]).
@@ 2.1 Path Types
In principle, given $a,b:A$, we could define the type of paths from $a$ to $b$ in $A$ as the type $\Sigma f:I\to A.(f~{}i_{0}=a)\times(f~{}i_{1}=b)$. However, experience with such a naïve formalization shows that it incurs a high number of laborious transportations along equalities that should be easy enough to infer automatically. Hence I instead follow the approach taken by Cubical Type Theory and related systems, and give an explicit axiomatization for path types, with corresponding rewrite rules to apply the associated equalities automatically:
@@
The introduction rule for path types corresponds to function abstraction
@@
We may then postulate the usual $\beta$-law as an identity for this type, along with special identities for application to $i_{0}$ and $i_{1}$. All of these are made into rewrite rules, allowing Agda to apply them automatically, and thus obviating the need for excessive use of transport: [^3] We could additionally postulate an $\eta$-law for path types, analogous to the usual $\eta$-law for function types; however, this is unnecessary for what follows, and so I omit this assumption.
@@
With this formalization of path types in hand, we can straightforwardly formalize the equivalent formulation of weak connectedness of $I$ given above. For this purpose, we first define the map idToPath that takes an identification $a≡b$ to a path from $a$ to $b$:
@@
A type $A$ is path-discrete if for all $a,b:A$ the map idToPath is an equivalence:
@@
which together imply that, if $A$ is discrete, then it is path-discrete:
@@
Although a full proof of canonicity is beyond the scope of this paper, I conjecture that adding these rules suffices to preserve canonicity, and I verify a few concrete cases of this conjecture later in the paper.
So much for the (weak) connectedness of $I$; let us now turn our attention to the other property we had previously stipulated of $I$, namely its strict bipointedness. As mentioned previously, we could simply postulate this stipulation directly as an axiom – however, for the purpose of proving parametricity theorems, a more prudent strategy is to instead formalize a class of $I$-indexed type families, whose computational behavior follows from this assumption (and which, in turn, implies it). Because these type families essentially correspond to the graphs of predicates and relations on arbitrary types, I refer to them as graph types.

### v2

@@ Abstract
Parametricity is a key metatheoretic property of type systems, which implies strong uniformity properties of the structure of types within systems possessing it. In recent years, various systems of dependent type theory have emerged with the aim of expressing such parametric reasoning in their internal logic. More recently still, such internal parametricity has been applied to solving various problems arising from the complexity of higher-dimensional coherence conditions in type theory. This paper presents a first step toward the unification, simplification, and extension of these methods for internalizing parametricity. Specifically, I argue that there is an essentially modal aspect of parametricity, intimately connected with the category-theoretic concept of cohesion. On this basis, I describe a general categorical semantics for modal parametricity, develop a corresponding framework of axioms (with computational interpretations) in dependent type theory, and demonstrate the practical utility of these axioms in managing the complexity of higher-dimension coherence by using them to painlessly derive induction principles for higher inductive types. In closing, I sketch the outlines of a more general synthetic theory of parametricity, with applications in domains ranging from homotopy type theory to the analysis of program modules.
@@
volume: NNjournal: Electronic Notes in Theoretical Informatics and Computer Sciencevolume: NN
@@
1 The Past, Present & Future of Internal Parametricity
@@
Reynolds [Rey83] began his seminal introduction of the concept of parametricity with the declaration that “type structure is a syntactic discipline for enforcing levels of abstraction.” In the ensuing decades, this idea of Reynolds’ has been overwhelmingly vindicated by the success of type systems in achieving modularity and abstraction in domains ranging from programming to interactive theorem proving. Yet much remains to be done in exploring the theoretical basis for and ramifications of this idea, and Reynolds’ particular solution to its formalization – the concept of relational parametricity.
@@
Reynolds’ original analysis of parametricity targeted the polymorphic $\lambda$-calculus (aka System F). Intuitively, polymorphic programs in System F cannot inspect the types over which they are defined and so must behave essentially the same for all types at which they are instantiated. To make this intuitive idea precise, Reynolds posed an ingenious solution in terms of logical relations, whereby every System F type is equipped with a suitable binary relation on its inhabitants, such that all definable terms in System F must preserve the relations defined on their component types. On this basis, Reynolds was able to establish many significant properties of the abstraction afforded by System F’s type structure, e.g. all closed terms of type $\forall X.X\to X$ are extensionally equivalent to the identity function.
Reynolds’ original analysis of parametricity for System F remained wholly external to System F itself. However, with the advent of systems of dependent type theory capable of conducting such analysis of proofs and programs internally, the possibility arises of finding an appropriate set of axioms for dependent type theory that would allow it to prove such parametricity theorems of its own type structure. Such internalization could moreover have far-reaching benefits for the overall theory of parametricity. Classical accounts of parametricity for various type systems have generally been “analytic” in nature – defined in terms of the particular syntactic constructs and models of the systems they target. On the other hand, axioms for parametricity internal to dependent type theory could provide a “synthetic” basis for parametricity whereby any system satisfying these axioms would necessarily enjoy parametricity and its consequences.
@@
The need for such an axiomatic specification of parametricity is made all the more apparent by recent developments in Homotopy Type Theory (HoTT) and related fields, wherein type theories with internal parametricity have been developed and applied by various authors to the solution of several open problems in these fields [CH21], [KS24]. Each such type theory has posed its own solution to the problem of internalizing parametricity, and although some commonalities exist between them, there is as yet no one framework that subsumes them all. In particular, the majority of these type theories have targeted only specific semantic models (usually some appropriately-chosen presheaf categories) rather than an axiomatically-defined class of models, and in this sense the analysis of parametricity offered by these type theories remains analytic, rather than synthetic. This limits any insight into whether and how these approaches to internal parametricity may be related to one another, generalized, or simplified.
@@
As a first step toward the unification, simplification, and generalization of the various approaches to parametricity, I propose an analysis of parametricity in terms of the category-theoretic concept of cohesion (and, more specifically, Lawvere’s concept of sufficient cohesion). Cohesion here refers to a particular relation between categories whereby a category is equipped with an adjoint string of modalities that together make its objects behave like spaces, whose points are bound together by an abstract property of cohesion, that constrains what maps may be defined between these spaces. Such constraint of the class of definable maps between objects is reminiscent of intuitive idea behind Reynolds’ original formulation of parametricity, particularly if one interprets the cohesion that binds types together as the structure of relations that inhere between them. Moreover, by inspection of the categorical models of extant type theories with internal parametricity, along with classical models of parametricity, one finds that many of these (if not quite all) exhibit cohesion, in this sense. The main contribution of this paper is thus twofold: 1) to show that this basic setup of cohesion is essentially all that is needed to recover classical parametricity results internally in dependent type theory, and 2) to show how this axiomatic framework can be applied in solving problems arising from the complexity of coherence conditions in HoTT – specifically, the problem of deriving induction principles for higher inductive types from their recursors. As a further illustration and verification of this idea, this paper is also a literate Agda document wherein the axioms for and theorems resulting from such parametricity via cohesion have been fully formalized and checked for validity. [^1] The full source of this file can be found at https://github.com/cbaberle/Parametricity-via-Cohesion.
{-# OPTIONS --rewriting --cohesion --flat-split --without-K #-}module parametricity-via-cohesion whereopen import Agda.Primitiveopen import Data.Emptyopen import Agda.Builtin.Unitopen import Agda.Builtin.Boolopen import Agda.Builtin.Sigmaopen import Agda.Builtin.Equalityopen import Agda.Builtin.Equality.Rewrite
@@ 2 Cohesion & Parametricity
where $\mathcal{E},\mathcal{S}$ are both topoi, $\Delta,\nabla$ are both fully faithful, and $\Pi$ preserves finite products. Given such an arrangement, we think of the objects of $\mathcal{E}$ as spaces and those of $\mathcal{S}$ as sets (even if $\mathcal{S}$ is not the category of sets), where $\Gamma$ is the functor that sends a space to its set of points, $\Delta$ sends a set to the corresponding discrete space, $\nabla$ sends a set to the corresponding codiscrete space, and $\Pi$ sends a space to its set of connected components. These in turn induce a string of adjoint modalities on $\mathcal{E}$:
@@
A concrete example of cohesion comes from the category of reflexive graphs $\mathbf{RGph}$, which is cohesive over the category of sets $\mathbf{Set}$ [Law05]. Here, $\Gamma$ is the functor that sends a reflexive graph to its set of vertices, $\Delta$ sends a set $V$ to the “discrete” reflexive graph on $V$ whose only edges are self-loops, $\nabla$ sends $V$ to the “codiscrete” (i.e. complete) graph where there is a unique edge between any pair of vertices in $V$, and $\Pi$ sends a reflexive graph to its set of (weakly) connected components. It is worth noting, at this point, that many classical models of parametricity (e.g. [Atk12]) are based upon semantic interpretation of type structure in terms of reflexive graphs. This, I wish to argue, is no accident, and the key property of reflexive graphs underlying such interpretations is precisely their cohesive structure. More generally, for any base topos $\mathcal{S}$, we may construct its corresponding topos $\mathbf{RGph}(\mathcal{S})$ of internal reflexive graphs, which will similarly be cohesive over $\mathcal{S}$. We can therefore in fact use the language of such internal reflexive graphs to derive parametricity results for any topos.
In fact, this same setup of cohesion is interpretable, mutatis mutandis, (c.f. [Sch13] §3.4) in the case where $\mathcal{E},\mathcal{S}$ are not (1-)topoi, but rather $\infty$-topoi, i.e. models of homotopy type theory [Shu18]. This allows us to use the language of homotopy type theory – suitably extended with constructs for the above-described modalities (the $\flat$ modality in particular, which, for technical reasons, cannot be axiomatized directly in ordinary HoTT) – to work synthetically with the structure of such a cohesive $\infty$-topos. For present purposes, we accomplish this by working in Agda with the --cohesion flag enabled, along with --without-K, which ensures compatibility with the treatment of propositional equality in HoTT. [^2] I also enable the --rewriting flag to allow specifying certain identities as rewrite rules in order to provide them with computational significance, and the --flat-split flag to allow pattern-matching over cohesive variables, which although not strictly necessary for the account of parametricity in what follows, is needed for some examples involving e.g. discreteness of the Booleans.
@@
    J⁻¹ :  {ℓ κ} {: Set ℓ} {: A}           (: (: A)  a ≡ b  Set κ)           {: A}  (: a ≡ b)  B b p  B a refl    J⁻¹ B refl b = b    transp⁻¹ :  {ℓ κ} {: Set ℓ} {a b : A}                (: A  Set κ)  a ≡ b  B b  B a    transp⁻¹ B p b = J⁻¹  a _  B a) p b
@@
The notion of contractibility expresses the idea that a type is essentially uniquely inhabited.
@@
The reader familiar with HoTT may note that, so far, I have not included anything relating to the univalence axiom – arguably the characteristic axiom of HoTT. In fact, this is by design, as a goal of the current formalization is to assume only axioms that can be given straightforward computational interpretations that preserve the property that every closed term of the ambient type theory evaluates to a canonical normal form (canonicity), so that these axioms give a constructive and computationally sound interpretation of parametricity. While the univalence axiom can be given a computational interpretation compatible with canonicity, as in Cubical Type Theory [CHS22], doing so is decidedly not a straightforward matter. Moreover, it turns out that the univalence axiom is largely unneeded in what follows, save for demonstrating admissibility of some additional axioms which permit more direct computational interpretations that are (conjecturally) compatible with canonicity. I thus shall have need for univalence only as a metatheoretic assumption. In this setting, univalence allows us to convert equivalences between types into identifications of those types, which may then be transported over accordingly.
@@
Of principal importance here is the $\flat$ modality, which (intuitively) takes a type $A$ to its corresponding discretization $\flat A$, wherein all cohesion between points as been eliminated. However, in order for this operation to be well-behaved, $A$ must not depend upon any variables whose types are not themselves discrete, in this sense. To solve this problem, the --cohesion flag introduces a new form of variable binding @♭ x : X, which judgmentally asserts that x is an element of the discretization of X, and enforces the restriction that X may only depend upon other variables bound with @♭. In this case, we say that x is a crisp element of X.
This notion in hand, we may define the $\flat$ modality as an operation on crisp types:
@@
This modality is then equipped with the following counit map $ε$:444arising from the fact that, semantically, $\flat$ is an idempotent comonad on crisp types.
    ε : {@♭ ℓ : Level} {@♭ A : Set ℓ}  ♭ A  A    ε (con x) = x
@@
Beyond such notions of discreteness, etc., what more is required for the sake of parametricity is a way of detecting when the elements of a given type are related, or somehow bound together, by the cohesive structure of that type.
@@
  • Strictly bipointed means that $I$ is equipped with two elements $i_{0},i_{1}:I$, such that the proposition $(i_{0}\equiv i_{1})\to\bot$ (i.e. $i_{0}\neq i_{1}$) holds in the internal language of $\mathcal{E}$.
@@
We could also, in principle, directly postulate the strict bipointedness of $I$, as an axiom having the form $i0≡i1\to⊥$. However, this is in fact unnecessary, as this axiom will instead follow from an equivalent formulation introduced in a following subsection.
@@
A type $A$ is weakly connected if and only if, for every discrete type $B$, any function $A\to B$ is essentially constant, in the sense of factoring through a contractible type.
@@
${{A}}$${{\smallint A}}$${{\flat B}}$${B}$$\scriptstyle{f_{\flat}}$$\scriptstyle{f^{\smallint}}$$\scriptstyle{\epsilon}$$\scriptstyle{\eta}$$\scriptstyle{f}$
@@
In order to express this property in Agda, it shall therefore be prudent first to introduce some machinery for ergonomically handling paths, analogous to the definition of path types in Cubical Type Theory (as in e.g. [CCHM16], [ABC+21]).
@@ 2.1 Path Types
In principle, given $a,b:A$, we could define the type of paths from $a$ to $b$ in $A$ as the type $\Sigma f:I\to A.(f~{}i_{0}=a)\times(f~{}i_{1}=b)$. However, experience with such a naïve formalization shows that it incurs a high cost of laborious transportations along identities that should be easy enough to infer automatically. Hence I instead follow the approach taken by Cubical Type Theory and related systems, and give an explicit axiomatization for path types, with corresponding rewrite rules to apply the associated equalities automatically:
@@
The introduction rule for path types corresponds to function abstraction.
@@
We may then postulate the usual $\beta$-law as an identity for this type, along with special identities for application to $i_{0}$ and $i_{1}$. All of these are made into rewrite rules, allowing Agda to apply them automatically, and thus obviating the need for excessive use of transport. [^5] We could additionally postulate an $\eta$-law for path types, analogous to the usual $\eta$-law for functions types; however, this is unnecessary for what follows, so I omit this assumption.
@@
With this formalization of path types in hand, we can straightforwardly formalize the equivalent formulation of weak connectedness of $I$ given above. For this purpose, we first define the map idToPath that takes an identification $a=b$ to a path from $a$ to $b$:
@@
A type $A$ is path-discrete if for all $a,b:A$, the map idToPath is an equivalence.
@@
which together imply that, if $A$ is discrete, then it is path-discrete.
@@
Although a full proof of canonicity is beyond the scope of this paper, I conjecture that adding these rules suffices to preserve canonicity.666perhaps not for Agda specifically, but for a more idealized system of type theory, such as Martin-Löf Type Theory I shall check a few specific cases of this conjecture later in the paper.
So much for the (weak) connectedness of $I$. Let us turn our attention now to the other property previously stipulated of $I$, namely strict bipointedness. As mentioned previously, we could simply postulate this stipulation directly as an axiom – however, for the purpose of proving parametricity theorems, a more prudent strategy is to instead formalize a class of $I$-indexed type families, whose computational behavior follows from this assumption (and which, in turn, implies it). Because these type families essentially correspond to the graphs of predicates and relations on arbitrary types, I refer to them as graph types.